OpenStudy (annie96):

Question 2.2. Solve the given system, using the substitution method. y = 3x – 7 6x – 2y = 12 (Points : 1) (14, 12) (12, 14) There is no solution. There are an infinite number of solutions. Question 3.3. Solve, using the substitution method. 5x + 2y = 7 y = x – 7 (Points : 1) (7, 0) (3, –4) (0, –7) (–1, –8)

OpenStudy (anonymous):

in 6x – 2y = 12 can you substitute y = 3x – 7 and find x ?

OpenStudy (annie96):

y=4x

OpenStudy (anonymous):

6x – 2y = 12 becomes 6x - 2(3x - 7) = 12 6x - 6x + 14 = 12 14 = 12 IMPOSSIBLE so there's no solution, right?

OpenStudy (annie96):

ok

OpenStudy (anonymous):

now can you do the same for the second one Q.3.3 ? Or, are you after the straight answer or want to learn how to do it ?

OpenStudy (annie96):

i want to learn how to do it.

OpenStudy (anonymous):

great!!! substitute y = x – 7 in 5x + 2y = 7 and u get 5x + 2(x-7) = 7 can u continue from here?

OpenStudy (annie96):

yeah

OpenStudy (annie96):

5x+2(-6)=7 5x-12=7 is that right.

OpenStudy (anonymous):

no dear ;) 2(x-7) = 2x - 14 can u continue?

OpenStudy (annie96):

yea

OpenStudy (annie96):

14x=-12y

OpenStudy (anonymous):

I'll write down 5x + 2(x-7) = 7 5x + 2x -14 = 7 , add +14 to both sides 5x + 2x - 14+14 = 7+14 , -14+14 = 0, 5x+2x = 7x, 7+14 = 21 7x = 21 , divide both sides by 7 7x/7 = 21/7 x = 3 now y = x - 7 - 3 - 7 = -4 so the answer is 3, -4 do u understand how we got it?

OpenStudy (annie96):

yeah thank you

OpenStudy (anonymous):

welcome I'd recommend u first try to strengthen ur skill on solving simple problems like 3(x + 11) = ? and -7(2 - x) = ?

OpenStudy (annie96):

ok